- Some work for Y arm alignment.
Transition from morning work (*:responsible person)
- We have maintenance today.
Report of work (*:responsible person)
- IOO, VAC: Pumping for MC failed yesterday. TP will start today.
- MC alignment can be started from this morning.
- CAL: Debugging ETMX Cal. Reached expected sensitivity. 60dB improvement was seen.
- Whitening for ETMY was installed yesterday and will be checked today.
- VIS: ITMY and BS TF are being measured yesterday and today.
- PR2 will be checked and tuned today.
- VIS: BS fishing rod is being fixed.
Work plan (*:responsible person)
- COM: Fiber noise measurement. Preparation of electronics for Y arm.
- AOS: Installation of WAB will start ETMX today.